OpenStudy (gabylovesyou):

Consider the line that passes through the points (2, 3) and (6, 11). Below are two different equations. Identify the true statement concerning both equations. Equation #1 Equation #2 y -11 = 2(x - 6) y - 3 = 2(x - 2) Only equation #1 represents the line that passes through the two given points. Only equation #2 represents the line that passes through the two given points. Both equation #1 and equation #2 represent the line that passes through the two given points. Neither equation #1 nor equation #2 represents the line that passes through the two given points.

OpenStudy (gabylovesyou):

@AccessDenied @Coolsector

OpenStudy (anonymous):

okay so we going to work together okay??

OpenStudy (gabylovesyou):

ok

OpenStudy (anonymous):

hold on working it out okay.. so what we suppose to do first??????????

OpenStudy (gabylovesyou):

solve each equation?

OpenStudy (anonymous):

okay got you..

OpenStudy (gabylovesyou):

right?? or wrong?

OpenStudy (anonymous):

oh okay;)

OpenStudy (anonymous):

The equations are in point slope form: y-y1=m(x-x1) where m=slope and (x1, y1) is whichever point you want.

OpenStudy (anonymous):

yes you you have to see what is right or wrong calebeavers each equation..

OpenStudy (anonymous):

Yep so if the slope is 2 then the answer would be Both equation #1 and equation #2 represent the line that passes through the two given points.

OpenStudy (gabylovesyou):

WHats the answer>

OpenStudy (anonymous):

okay calebeavers gave you the answer..

OpenStudy (gabylovesyou):

ohh lol thanks :)

OpenStudy (anonymous):

anytime ...